OpenStudy (anonymous):

how to solve this limit stepwise? lim(x-->infinity) 2e^x+3e^-x/5e^x-7e^-x

OpenStudy (turingtest):

multiply by\[{e^{-x}\over e^{-x}}\]then evaluate

OpenStudy (anonymous):

I don't get it

OpenStudy (turingtest):

what do you get after that multiplication?

OpenStudy (anonymous):

2e^x(e^-x)+4e^-x/5e^x-8e^-x

OpenStudy (turingtest):

no, that is not\[{2e^x+4e^{-x}\over5e^x-8e^{-x}}\cdot{e^{-x}\over e^{-x}}\]check your algebra and also, how do you simplify \(e^x\cdot e^{-x}\) ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

so it's 2+4e^-x/5-8e^-x which gives u 2/5?

OpenStudy (turingtest):

yes :)

OpenStudy (anonymous):

thank you!!!!

OpenStudy (turingtest):

welcome :)

OpenStudy (anonymous):

i suck at limits :(

OpenStudy (turingtest):

btw it should be \(e^{-2x}\) after multiplication, but that didn't matter since those terms vanish anyway limits' like all things, come with practice once you've seen all the tricks, you'll get used to it :)

OpenStudy (turingtest):

^ the \(4e^{-2x}\) part I meant (since \(e^{-x}\cdot e^{-x}=e^{-2x}\))

OpenStudy (anonymous):

thank you! :)